OpenStudy (anonymous):

Determine how much time is required for an investment to double in value if interest is earned at a rate of 6.25% compounded continuously ?

OpenStudy (anonymous):

what is continuously?

OpenStudy (anonymous):

Hmmm I would say every single day so its compounded 365 days But lemme just ask inquire with @dumbcow Either way there is an equation to determine the Effective Annual Rate when its compounded continuously aka daily 365 days a year The formula is as follows \[e^{r}-1=e^{.0625}-1=.0645\] This is equivalent to 6.45%

OpenStudy (anonymous):

So assuming that we have an EAR of 6.45 % we can have the following formula Lets say our investment is "a" We start off with "a" and we are wondering how long it will take till our investment doubles to 2a Lets use our regular folrmula \[2a=a(1+.645)^x\] Solve for x Do you follow what I have done??? Any questions?

OpenStudy (dumbcow):

@wahahaha , compound continuously is even more than 365 days, its like every second but just use the "e" and you are fine to double you want growth to be factor of 2 \[e^{.0625 t} = 2\] \[t = \frac{\ln 2}{.0625}\]

OpenStudy (anonymous):

@dumbcow Yaaaa okkk Its just a branch off of what I did above but just in a more efficient manner K thanks

OpenStudy (dumbcow):

yeah you are not wrong, but you will have a rounding error

OpenStudy (anonymous):

Ok I gotcha ... Thanks

OpenStudy (anonymous):

thanks to you both!